OpenStudy (anonymous):

f(x)=√1-x, g(x)=1/x find (f of g)(10)

OpenStudy (whpalmer4):

\[f(x)=\sqrt{1-x}\]\[g(x)=\frac{1}{x}\]Evaluate f(g(10)): \[g(10)=\frac{1}{10}\]\[f(g(10))=f(\frac{1}{10})=\sqrt{1-\frac{1}{10}}=\sqrt{\frac{10}{10}-\frac{1}{10}}=\frac{3}{\sqrt{10}}\]

OpenStudy (whpalmer4):

Note that you could also have simplified the expression for f(g(x)) before plugging in any numbers: \[f(g(x))=\sqrt{1-g(x)} = \sqrt{1-\frac{1}{x}}\]